This thread has been locked.
If you have a related question, please click the "Ask a related question" button in the top right corner. The newly created question will be automatically linked to this question.
LM5002: LM5002 SEPIC with coupled inductors
Part Number: LM5002
Hi E2E forum!
We've used LM5002 in SEPIC configuration as per example below:
We have just changed the Feedback resistors to have 15V as output.Our specification is Vin=14V to 20V, Vout=15V at 50mA
From the operation side the SEPIC is working quite good: sweeping from 14V to 20 the output is always stable at 15V. However something strange on the output is happening, it seems ripple but in fact seems to be more a switching noise issue.At the beginning we was using as switching Diode (D2) a Vishay V8PAL45 but seems to have a high junction capacitance(1400pF) so we changed with a SS16 (less than 100pF).Changing diode and output capacitor (from electrolitic to ceramic) have improved the ripple as below image to 200mV.
Zoom on the peak-pulse:
All the test has been performed with a 50mA load.We're basically looking for suggestion on hot to improve the AC-Ripple/Switching noise at the output: We tried to add a RC snubber in parallel to D2 but unfortunately this have not helped.Anyone has got similar issues?
Thanks in advance
I done some further investigation, and yes seems really to be more a Switching noise instead of AC Ripple:
Measurement has been made with the tip-ring to shorten as soon as possible the ground path.
Trace #1 (Yellow): Ouput , coupling in AC
Trace#2 (Green): Pin N° 1 LM5002 (coupled in DC).
We are glad that we were able to resolve this issue, and will now proceed to close this thread.
If you have further questions related to this thread, you may click "Ask a related question" below. The newly created question will be automatically linked to this question.
In reply to Davide Dal Fra:
I think the behavior is mainly caused by the probe setup as show in the image
Best RegardsJasper LiBoost Converter Solution
In reply to Jasper Li:
Thankyou for your reply.However, the measurement was made correctly by using the tip-spring tool on the oscilloscope probe instead of using the normal gnd wire+clamp.
Yes exactly the waveform has been measured directly on the output capacitor.Adding a 0603 100nF capacitor (we have finished 0402) have reduced about 5/7% the spikes.
We'll try add a smaller LC filter, and i'll let you know.
Thanks in advance for your support
All content and materials on this site are provided "as is". TI and its respective suppliers and providers of content make no representations about the suitability of these materials for any purpose and disclaim all warranties and conditions with regard to these materials, including but not limited to all implied warranties and conditions of merchantability, fitness for a particular purpose, title and non-infringement of any third party intellectual property right. No license, either express or implied, by estoppel or otherwise, is granted by TI. Use of the information on this site may require a license from a third party, or a license from TI.
TI is a global semiconductor design and manufacturing company. Innovate with 100,000+ analog ICs andembedded processors, along with software, tools and the industry’s largest sales/support staff.